# PI3D: Efficient Text-to-3D Generation with Pseudo-Image Diffusion

### 14 Dec 2023 v1

@@ 1 Introduction
In order to generate 3D assets in scenarios where 3D priors are difficult to obtain, researchers have recently been keen to utilize various 3D representations [24, 41, 21, 35] to lift 2D diffusion models with Score Distillation Sampling (SDS) [27, 40]. However, applying only 2D prior to optimization is prone to hinder the 3D nature, inevitably compromising 3D fidelity and consistency. 3D knowledge is a key issue that cannot be bypassed for high-quality 3D generation.
@@ 2.1 Diffusion-based 3D Generative Model
This line of research focuses on training 3D diffusion models on popular 3D datasets [5, 9, 8], each with distinct representations of 3D objects. As these models have access to 3D raw data, they inherently acquire 3D knowledge inductively from the dataset. Point·E [25] trains two point-cloud diffusion models conditioned on the text prompt or single-view image to generate a low-resolution point cloud and up-sample it respectively. It can further obtain meshes from the point cloud via SDF prediction. Shap·E [19] trains the diffusion model on the 3D implicit function. It is capable of simultaneously producing NeRF and SDF with two heads, but it struggles with semantics when faced with complex prompts. NFD [37], 3DGen [11], and DiffTF [3] adopt diffusion on the triplane representation to achieve highly efficient mesh generation but they constrain the generation to be category-specific on ShapeNet [5] and OmniObject3D [43]. All of these models actually consider only 3D data, which leads to their poor ability to generate reasonable objects when confronted with some complex concepts.
@@
This line of work is typically focused on optimizing a 3D representation based on a pre-trained 2D diffusion model. The foundational technique involved is Score Distillation Sampling (SDS), first proposed in [27, 40]. SDS aims to align the rendering results of a parameterized renderer with the distribution described by a pre-trained image diffusion model, achieving parameter space sampling. In DreamFusion [27] and SJC [40], 3D knowledge originates solely from spatial representations and the view-dependency prompting, which are relatively weak. This limitation increases the likelihood of 3D inconsistency, such as the Janus problem, since there is no direct informative constraint between different views.
Perp-Neg [2] mitigates this problem by correcting the direction of parameter updates guided by the 2D diffusion model but leaves the lack of a 3D prior unsolved. Zero-1-to-3 [22] proposes to train a 2D viewpoint transformation model conditioned on the input view and the relative camera transformation, injecting implicit 3D knowledge into the 2D diffusion model. The viewpoint-conditioned diffusion model can be further used for generating 3D consistent objects from a single view via SDS. Magic123 [28] fuses and balances guidance from pre-trained 2D diffusion model [31] and viewpoint transformation model [22], achieving detailed image-to-3D results. MVDream [36] models the distribution of multi-view images and thus greatly improves the stability of existing 2D-lifting methods. Although methods in this section are usually capable of generating an attractive appearance, they all require tons of optimizing iterations and are of lower diversity.
3 Preliminary
Diffusion models [38, 39, 15] are a class of generative models capable of generating diverse samples from pure noise by progressively denoising it. Given a sample $z_{0}$ from the real data distribution, a forward diffusion process is defined as gradually adding noise to $z_{0}$ to form a sequence $\{z_{t}=\alpha_{t}z_{0}+\sigma_{t}\epsilon|t=1,2,...,T\}$, where $\epsilon\sim\mathcal{N}(0,1)$ is a gaussian noise. Here, $\alpha_{t},\sigma_{t}$ are pre-defined coefficients ensuring that $z_{t}$ is close to the real sample at the start and becomes Gaussian noise towards the end of the forward process, respectively. To generate samples, a diffusion model $\epsilon_{\phi}(\cdot;t)$ is trained to approximate the reverse process and gradually denoise the sample by predicting the noise level at the timestep $t$. A common training objective is a weighted error prediction MSE loss:
$\mathcal{L}=\mathbb{E}_{t\sim[1,...,T],\epsilon\sim\mathcal{N}(0,1)}[w(t)||\epsilon_{\phi}(\alpha_{t}z_{0}+\sigma_{t}\epsilon;t)-\epsilon||_{2}]$ (1)
@@ 4 Method
[…]
Figure 3: The illustration of our depth-aware triplane fitting.
@@ 4.1 Depth-Aware Triplane Fitting
$\text{where}\ T_{i}=\exp(-\sum_{j=1}^{i-1}\sigma\delta(p_{j}))$
$\{p_{i}\}_{i=1}^{N}$ denotes the sampled points along that ray. We do not involve additional positional embedding like in vanilla NeRF [24], since we want the decoder could be position-agnostic, ensuring that what is similar on the rendered images is also similar on the triplane. To encourage a correct underlying geometry, we also supervise the object mask $\hat{M}(r)$ and the expected depth $\hat{D}(r)$ along that ray, which are given by:
@@
The overall loss on rendering views is as follows:
@@
$\mathcal{L}_{\text{mask}}$ $=\lambda_{\text{mask}}\sum_{r}\text{BCE}(\hat{M}(r),M(r))$
$\mathcal{L_{\text{depth}}}$ $=\lambda_{\text{depth}}\sum_{r}(M(r)\otimes||\hat{D}(r)-D(r)||_{1})$
@@
In addition to encouraging the triplane to fit the object well, we still want it to be friendly to the later diffusion process. So it is important to put several regularizations on the triplane itself. Inspired by [3], we use L2 and TV loss as follows:
@@
In order for the geometry to converge fast, we render 3 axis-aligned views in orthogonal projection and adopt these masks $O_{xy},O_{yz},O_{xz}$ on 3 planes to constraint where it has non-zero features, which is basically equivalent to a 3-view visual hull constraint.
@@ 4.3 Lightweight Refinement
Upon obtaining the diffusion model, two intuitive ways to generate 3D models from input texts are considered. One can directly sample new 3D models within seconds from it and further refine it with 2D SDS. Alternatively, we can view the 3D diffusion model as a prior and jointly optimize a triplane-based NeRF with 2D and 3D SDS, like Magic123 [28] to guarantee the 3D consistency. In practice, we find that sampling a 3D model first using our method and then refining it is a simple yet effective way to produce high-quality objects and guarantee 3D consistency. We continue to optimize the triplane generated by the pseudo-image diffusion along with the shared decoder with SDS loss. To prevent the 3D shape from being over-optimized, we refine the appearance and geometry by sampling only small noise levels in SDS. Thanks to a good initialization, it significantly accelerates the convergence compared to previous optimization-based methods [27, 6] and thus arrives at efficient 3D content generation.
@@ 5.1 Implementation Details
We defoine the triplane size as $H=W=256,C=6$. The shared decoder used in volume rendering is a 4-layer MLP, with each hidden layer having a width of 32. It is trained on a set of 500 objects. Subsequently, we freeze the shared decoder and independently fit the triplane representation for each object, implementing a linearly decayed learning rate. The balancing coefficients for loss items are as follows: $\lambda_{\text{L1}}=0.2,\lambda_{\text{mask}}=0.1,\lambda_{\text{depth}}=0.5,\lambda_{\text{L2}}=0.05,\lambda_{\text{TV}}=0.05,\lambda_{\text{hull}}=1.0$. The triplanes are scaled to $[-1,1]$ to ensure a similar variance with images for easy diffusion training later. The process of fitting one object on a single A100 GPU takes 2 minutes.
@@
In pseudo-image sampling, we use DDPM [15] scheduler for 50 sample steps with a classifier free guidance [14] scale of 5.0. It takes 3 seconds to sample a triplane with the $256\times 256$ resolution. We implement the subsequent refinement based on threestudio [10], utilizing DeepFloyd-IF [1] as the 2D guidance model. The triplane-based NeRF is optimized for 2000 steps with a batch size of 1, requiring 3 minutes on an A100 GPU. The timestep $t$ is sampled in $\mathcal{U}(0.1,0.5)$ to introduce small corrections and prevent excessive changes. The classifier-free guidance adopted in SDS is 20.
@@ 5.2 Comparisons with Other Text-to-3D Methods
We compare PI3D with general text-to-3D diffusion models, Point·E [25] and Shap·E [19] and SDS-based method DreamFusion [27]. To evaluate the fidelity, we adhere to the metrics used in prior works [19, 27], namely CLIP R-Precision and CLIP similarity score [13]. CLIP R-precision measures the accuracy of retrieving text prompts using rendered images based on CLIP similarity. We select 4 views at different camera poses with azimuth angles $0,90,180,270$ degrees and the same elevation angles $15$ degrees, performing multiple retrievals to obtain an average. We compare these metrics on the 415 prompts from the DreamFusion gallery.
We first compare the pseudo-image diffusion model with the baseline text-to-3D diffusion models. The quantitative results are reported in Tab. 1. Our model significantly outperforms the other two methods in all the metrics, revealing that it could generate more text-aligned 3D models. We also demonstrate the generation quality through some qualitative cases (see Fig. 4), each presented with 3 rendered views. The baselines, Point·E [25] and Shap·E [19], can generate simple objects that are part of complex prompts but struggle to capture the semantics of intricate text prompts, for example, depicting ”corn fields” without the tractor or a ”penguin” without the skiing activity. We believe the limited generalization ability is partly because they are both trained only on 3D datasets with variable quality of text annotation and 3D geometry. However, leveraging the similarity between pseudo-images and 2D real images, our method can learn rich concepts from limited 3D data and achieve reasonable generation quality even with prompts involving concept combination.
We also compare the refinement stage with DreamFusion [27] on the generation efficiency and quality. We present the implementation for DreamFusion-IF in threestudio [10]. We compare their efficiency and quality in Tab. 2 and Fig. 5, respectively. The baseline method requires tons of iterations in hours while still performing poorly in 3D consistency and convergence. Our method can generate 3D consistent models and converge much faster with an even smaller batch size in only 3 minutes by leveraging a good 3D initialization sampled from the diffusion model.
Table 2: Comparisons of the computation costs of DreamFusion-IF [27] and PI3D. Time costs are evaluated on a single A100 GPU. We achieve a 30$\times$ speedup.
Method Iterations Time Batch size
DreamFusion-IF [27] 15000 $\sim$ 1.5 hr 4
Ours 2000 3 min 1
@@ 6 Limitations
Although PI3D demonstrates the ability to efficiently generate 3D consistent shapes from text prompts, there still exist some limitations left for the future to address. First, our training procedure involves a per-object triplane fitting phase, which leads to an exponential increase in training costs when scaling up to larger 3D datasets. This might be solved by encoding triplanes from point clouds or multi-view images with an autoencoder architecture like [25, 11]. Additionally, the capacity of our scene representation is limited by the feature dimensions. When it comes to more detailed 3D generation, higher dimensional features are inevitably required, which raises the number of pseudo-images and poses a challenge in modeling the relationships between them. This may be solved by designing more explicit and effective feature interactions.

### 21 Apr 2024 v2

@@ 1 Introduction
In order to generate 3D assets in scenarios where 3D priors are difficult to obtain, researchers have recently been keen to utilize various 3D representations [31, 49, 24, 43] to lift 2D diffusion models with Score Distillation Sampling (SDS) [34, 48]. However, applying only 2D prior is prone to hinder the 3D nature, inevitably compromising 3D fidelity and consistency. 3D knowledge is a key issue that cannot be bypassed for high-quality 3D generation.
@@ 2.1 Diffusion-based 3D Generative Model
This line of research focuses on training 3D diffusion models on popular 3D datasets [5, 9, 8], each with distinct representations of 3D objects. As most of these methods have access to 3D raw data, they inherently acquire 3D knowledge inductively from the dataset. Point·E [32] trains two point-cloud diffusion models conditioned on the text prompt or single-view image to generate a low-resolution point cloud and up-sample it respectively. It can further obtain meshes from the point cloud via SDF prediction. Shap·E [20] trains the diffusion model on the 3D implicit function. It is capable of simultaneously producing NeRF and SDF with two heads, but it struggles with semantics when faced with complex prompts. NFD [45], 3DGen [11], and DiffTF [3] adopt diffusion on the triplane representation to achieve highly efficient mesh generation but they constrain the generation to be category-specific on ShapeNet [5] and OmniObject3D [52]. Some other methods [23, 22] build 3D diffusion models with only posed 2D images on smaller datasets [38]. All these methods actually utilize only poorly labeled 3D data or limited image data, which leads to their poor ability to generate reasonable objects when confronted with some complex concepts.
@@
Several recent works focus on optimizing a 3D representation by leveraging a pre-trained 2D diffusion model. The foundational technique involved is Score Distillation Sampling (SDS) [34, 48]. SDS aims to align the rendering results of a parameterized renderer with the distribution described by a pre-trained image diffusion model, achieving parameter space sampling. In DreamFusion [34] and SJC [48], 3D knowledge originates solely from spatial representations and the view-dependency prompting, which are relatively weak. This limitation increases the likelihood of 3D inconsistency, such as the Janus problem, since there is no direct informative constraint between different views. ProlificDreamer [50] and HiFA [53] reduce the oversaturation and greatly increase the generation quality by reinventing the optimization target. Perp-Neg [2] mitigates the inconsistency problem by correcting the direction of parameter updates guided by the 2D diffusion model but leaves the lack of a 3D prior unsolved. Zero-1-to-3 [27] proposes to train a 2D viewpoint transformation model conditioned on the input view and the relative camera transformation, injecting implicit 3D knowledge into the 2D diffusion model. Magic123 [35] fuses and balances guidance from pre-trained 2D diffusion model [39] and viewpoint transformation model [27], achieving detailed image-to-3D results. MVDream [44] models the distribution of multi-view images and thus greatly improves the stability of existing 2D-lifting methods. ATT3D [29] amortizes the optimization cost among prompts by introducing a prompt-conditioned parameter prediction hyper-network. DreamGaussian [46] accelerates the generation by utilizing real-time rendered 3D representation. Although these methods are usually capable of generating an attractive appearance, they all require tons of optimizing iterations and are of lower diversity.
Some other methods view 3D generation as multi-view generation combined with reconstruction [25, 28, 47, 26]. The reconstruction process can be optionally replaced by a feed-forward model [26, 18]. These approaches allow for fast 3D generation but are often jointly affected by minor inconsistencies of multi-view images and errors in the reconstruction process.
[…]
Figure 3: We fit the triplane representation for each object with the supervision of the rendered RGB images, binary mask, and object depth. We also adopt axis-aligned masks on the triplane to encourage the geometry to approach the correct surface faster.
@@ 3.1 Depth-Aware Triplane Fitting
$\text{where}\ T_{i}=\exp(-\sum_{j=1}^{i-1}\delta\sigma(p_{j}))$
$\{p_{i}\}_{i=1}^{N}$ denotes the sampled points along that ray. We do not involve additional positional embedding like in vanilla NeRF [31], since we want the decoder to be position-agnostic, ensuring that what is similar on the rendered images is also similar on the triplane. To encourage a correct underlying geometry, we also supervise the object mask $\hat{M}(r)$ and the expected depth $\hat{D}(r)$ along that ray, which are given by:
@@
The losses on rendering views are as follows:
@@
$\mathcal{L}_{\text{mask}}=\sum_{r}\text{BCE}(\hat{M}(r),M(r))$ (7)
$\mathcal{L_{\text{depth}}}=\sum_{r}(M(r)\otimes||\hat{D}(r)-D(r)||_{1})$ (8)
@@
In addition to encouraging the triplane to fit the object well, we still want it to be friendly to the later diffusion process. So we put several regularizations on the triplane itself. Inspired by [3], we use L2 and TV loss as follows:
@@
In order for the geometry to converge fast, we render 3 axis-aligned views in orthogonal projection and adopt these masks $O_{xy},O_{yz},O_{xz}$ on 3 planes to constrain where it has non-zero features, which is basically equivalent to a 3-view visual hull constraint.
@@ 3.3 Lightweight Refinement
Upon obtaining the diffusion model, two intuitive ways to generate 3D models from input texts are considered. One can directly sample new 3D models within seconds from it and further refine it with 2D SDS. Alternatively, we can view the 3D diffusion model as a prior and jointly optimize a triplane-based NeRF with 2D and 3D SDS, like Magic123 [35] to guarantee the 3D consistency. In practice, we find that sampling a 3D model first using our method and then refining it with only 2D diffusion models is a simple yet effective way to produce high-quality objects and guarantee 3D consistency. In the refinement, we continue to optimize the triplane generated by the pseudo-image diffusion along with the shared decoder with SDS loss. To prevent the 3D shape from being over-optimized, we refine the appearance and geometry by sampling only small noise levels in SDS. A good coarse 3D model provided by the pseudo-image diffusion gives 2D SDS a strong appearance prior, making it converge much faster than starting from scratch [34, 6, 50, 44].
@@ 4.1 Implementation Details
We define the triplane size as $H=W=256,C=6$. The shared decoder used in volume rendering is a 4-layer MLP, with each hidden layer having a width of 32. It is trained on a set of 500 objects. Subsequently, we freeze the shared decoder and independently fit the triplane representation for each object, implementing a linearly decayed learning rate. The balancing coefficients for loss items are as follows: $\lambda_{\text{L1}}=0.2,\lambda_{\text{mask}}=0.1,\lambda_{\text{depth}}=0.5,\lambda_{\text{L2}}=0.05,\lambda_{\text{TV}}=0.05,\lambda_{\text{hull}}=1.0$. The triplanes are scaled to $[-1,1]$ to ensure a similar variance with images for easy diffusion training later. The process of fitting one object on a single A100 GPU takes 2 minutes.
@@
In pseudo-image sampling, we use DDPM [15] scheduler for 50 sample steps with a classifier free guidance [14] scale of 5.0. It takes 3 seconds to sample a triplane with the $256\times 256$ resolution. We implement the subsequent refinement based on threestudio [10], utilizing DeepFloyd-IF [1] as the 2D guidance model. The triplane-based NeRF is optimized for 2000 steps with a batch size of 1, typically requiring 3 minutes on an A100 GPU. The timestep $t$ is sampled in $\mathcal{U}(0.1,0.5)$ to introduce small corrections and prevent excessive changes. The classifier-free guidance adopted in SDS is 20.
@@
Figure 6: We replace the 2D guidance used in the refinement with those in other methods. We can achieve comparable or better results on significantly fewer iterations.
@@ 4.2 Comparisons with Other Text-to-3D Methods
We compare PI3D with general text-to-3D diffusion models, Point·E [32] and Shap·E [20] and SDS-based method DreamFusion [34], Fantasia3D [6], ProlificDreamer [50], MVDream [44]. To evaluate the fidelity, we adhere to the metrics used in prior works [20, 34], namely CLIP R-Precision and CLIP similarity score [13]. CLIP R-precision measures the accuracy of retrieving text prompts using rendered images based on CLIP similarity. We select 4 views at different camera poses with azimuth angles ($0,90,180,270$ degrees) and the same elevation angle ($15$ degrees), performing multiple retrievals to average a score. All 415 testing prompts are from the DreamFusion gallery.
We first compare the pseudo-image diffusion model with the baseline text-to-3D diffusion models. The quantitative results are reported in Tab. 1. Our model significantly outperforms the other two methods in all the metrics, revealing that it could generate more text-aligned 3D models. We also demonstrate the generation quality through some qualitative cases (see Fig. 4), each presented with 3 rendered views. The baselines, Point·E [32] and Shap·E [20], can generate simple objects that are part of complex prompts but struggle to capture the semantics of intricate text prompts, for example, depicting “corn fields" without the tractor or a “penguin" without the skiing activity. We believe the limited generalization ability is partly because they are both trained only on 3D datasets with variable quality of text annotation and 3D geometry. However, leveraging the similarity between pseudo-images and 2D real images, our method can learn rich concepts from limited 3D data and achieve reasonable generation quality even with prompts involving concept combination.
We also compare the refinement stage with SDS-based methods on the generation efficiency and quality. Several examples by DreamFusion-IF from threestudio [10] and our method are compared in Fig. 5. The baseline method requires tons of iterations in hours while still performing poorly in 3D consistency and convergence. Our method can generate 3D consistent models and converge much faster with an even smaller batch size in only 3 minutes by leveraging a good 3D initialization sampled from the diffusion model. Also, our method is compatible with other SDS variants. As demonstrated in Fig. 6, we compare our method with other SDS-based methods by switching the 2D guidance to the corresponding variants. We can achieve superior or comparable quality with even much faster convergence. In terms of efficient generation, ATT3D [29] has an inference runtime of 1s, much faster compared to optimization-based methods. But we contend that it is much less generalizable than our method. Limited by a prompt set, it can only generate new shapes by combining seen words while PI3D supports complex concepts. Our sampling takes around 3s and ATT3D has a runtime of 1s. As it is not open-sourced, we show the results from their paper in Fig. 7. Furthermore, it employs a deterministic mapping network with no diversity.
Figure 7: Our diffusion model beats ATT3D in terms of 3D consistency and can produce diverse results.
@@ 5 Limitations
Although PI3D demonstrates the ability to efficiently generate 3D consistent shapes from text prompts, there still exist some limitations left for the future to address. First, our training procedure involves a per-object triplane fitting phase, which leads to a linear increase in training costs when scaling up to larger 3D datasets. This might be solved by encoding triplanes from point clouds or multi-view images with an autoencoder architecture like [32, 11]. Additionally, the capacity of our scene representation is limited by the feature dimensions. When it comes to more detailed 3D generation, higher dimensional features are inevitably required, which raises the number of pseudo-images and poses a challenge in modeling the relationships between them. This may be solved by designing more explicit and effective feature interactions.
@@ 6 Conclusion
Acknowledgement This work was supported by the National Key Research and Development Program of China (No. 2023YFF0905104), the Natural Science Foundation of China (No. 62132012, No. 62361146854), Beijing Municipal Science and Technology Project (No. Z221100007722001) and Tsinghua-Tencent Joint Laboratory for Internet Innovation Technology.